OpenStudy (anonymous):

WHat is the diameter of a circle w/ a circumference that is 443 feet

OpenStudy (anonymous):

Do you know the formula for circumference?

OpenStudy (anonymous):

no

OpenStudy (anonymous):

\[C=2\pi^2\] Since we are given the circumference, we can plug that into the equation.

OpenStudy (studygurl14):

Circumference = diameter times pi

OpenStudy (studygurl14):

Hey @JackSmiley I think you made a mistake...

OpenStudy (anonymous):

My apologies, I meant, \[2\pi(r^2)\]

OpenStudy (anonymous):

okay what do i do

OpenStudy (anonymous):

Since we know C, the equation would be: \[443=2\pi(r^2)\] From this, we can solve for r.

OpenStudy (anonymous):

so 443/2

OpenStudy (anonymous):

Yes, \[443/2=\pi(r^2)\] Do you know what to do after this?

OpenStudy (anonymous):

um is pi 3.14

OpenStudy (anonymous):

Yes. :) But most scientific calculators have a pi button, and it is better to use that instead of an estimate. :)

OpenStudy (anonymous):

Wait i don't get that part the last part

OpenStudy (anonymous):

Which part?

OpenStudy (anonymous):

the pi(radius2)

OpenStudy (anonymous):

Pi is a number, as 2 or any other integer, which means you can divide it as you would any other number, just as you did with the 2.

OpenStudy (anonymous):

but how do i find radius

OpenStudy (anonymous):

Radius squared, or r^2, can be solved for by taking the square root fo both sides, leavig you with \[r^2=(term)\] \[\sqrt{(term)}=r\]

OpenStudy (anonymous):

wait how do i find radius though

OpenStudy (anonymous):

Radius = r

OpenStudy (anonymous):

\[443/2=\pi(r^2)\]\[443/(2\pi)=r^2\]\[r=\sqrt{443/(2\pi)}\]

OpenStudy (studygurl14):

hey @JackSmiley, you still got the formula wrong...